find aagles angle A=5x+8
angle B=x+4

Answer by
rapaljer(4551)

5x+8 +x+4 = 90
6x+12 =90
6x=78
x=13

Angle A = 5x +8
Angle A = 5*13 + 8 = 65+8=73 degrees

Angle B = x+4
Angle B = 13+4 = 17 degrees

Check: 73 + 17=90

R^2

Dr. Robert J. Rapalje
Altamonte Campus
Seminole State College
Florida